# Noise estimation of SDE from a single data trajectory

### v1

@@ 3 Main Results
3.3 Proof of Step 4.
Proof. Applying Girsanov theorem, plugging the dynamic $dB_{t}^{\mathbb{P}}=dB_{t}^{\mathbb{Q}}-\frac{\mu(X_{t})}{\sigma(X_{t})}dt$ into the expanison (32) from Step 3, we have
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu$
$\qquad+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du+\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}dr\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
$\qquad+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
@@
$\qquad+\left((\tilde{\mu}^{\prime}(X_{s}))^{2}\sigma(X_{s})+\tilde{\mu}^{\prime\prime}(X_{s})\tilde{\mu}(X_{s})\sigma(X_{s})\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{v}^{\mathbb{Q}}-\frac{\mu(X_{v})}{\sigma(X_{v})}dv\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}drdu$
$\qquad+\left(\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})+\tilde{\mu}^{\prime\prime}(X_{s})\tilde{\mu}(X_{s})\sigma(X_{s})\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}dv\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad+\left(\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})+\sigma^{\prime\prime}(X_{s})(\tilde{\mu}(X_{s}))^{2}\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}dvdr\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
$\qquad+\left(\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\sigma(X_{s})+\tilde{\mu}^{\prime\prime}(X_{s})(\sigma(X_{s}))^{2}\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{v}^{\mathbb{Q}}-\frac{\mu(X_{v})}{\sigma(X_{v})}dv\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad+\left(\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\sigma(X_{s})+\sigma^{\prime\prime}(X_{s})\sigma(X_{s})\tilde{\mu}(X_{s})\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{v}^{\mathbb{Q}}-\frac{\mu(X_{v})}{\sigma(X_{v})}dv\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dr\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
$\qquad+\left((\sigma^{\prime}(X_{s}))^{2}\tilde{\mu}(X_{s})+\sigma^{\prime\prime}(X_{s})\sigma(X_{s})\tilde{\mu}(X_{s})\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}dv\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
@@
$\qquad\qquad\qquad\qquad\times\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{v}^{\mathbb{Q}}-\frac{\mu(X_{v})}{\sigma(X_{v})}dv\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
@@
$X_{t}$ $=X_{s}-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
$\qquad+\left(\mu^{\prime}(X_{s})\sigma(X_{s})-\mu(X_{s})\sigma^{\prime}(X_{s})-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})^{2}-\frac{1}{2}\sigma(X_{s})^{2}\sigma^{\prime\prime}(X_{s})\right)\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{r}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})^{2}\int_{0}^{t}\int_{0}^{s}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}du\circ dB_{s}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{r}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
$+\left((\sigma^{\prime}(X_{s}))^{2}\sigma(X_{s})+\sigma^{\prime\prime}(X_{s})(\sigma(X_{s}))^{2}\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{v}^{\mathbb{Q}}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\varepsilon^{\tilde{\mu},\sigma}+\varepsilon^{\mu/\sigma}.$
@@
3.4 Proof of Step 5
Proof. Comparing the coefficients of (20) from Step 2 and (23) from Step 4, for the expansion of $X_{t}$ at the initial point $X_{s}$ on interval $[s,t)$, we have the following equalities,
@@ 3.4 Proof of Step 5
Following the above observation, the parameter $\psi^{t_{i},t_{i+1}}_{21}$ can be found corresponding to the initial value $X_{t_{i}},\text{ for }i=0,1,2,\cdots,n$. With this, the equation (35) holds true for each $t_{i},i=0,1,2,\dots,n$. For example, Eq. (35) equivalently written as
$\mu^{\prime}(X_{t_{i}})\sigma(X_{t_{i}})-\mu(X_{t_{i}})\sigma^{\prime}(X_{t_{i}})-\frac{1}{2}\sigma(X_{t_{i}})\sigma^{\prime}(X_{t_{i}})^{2}-\frac{1}{2}\sigma(X_{t_{i}})^{2}\sigma^{\prime\prime}(X_{t_{i}})=\psi_{21}^{t_{i},t_{i+1}},\quad\forall\;i=0,1,2,\dots,n$ (36)
This equation can be written compactly as
$\mu^{\prime}(X_{r})=a(X_{r})\mu(X_{r})+b(X_{r}),$ (37)

### v2

@@ Appendix B Proof of Step 3
Appendix C Proof of Step 4
Applying Girsanov theorem, plugging the dynamic $dB_{t}^{\mathbb{P}}=dB_{t}^{\mathbb{Q}}-\frac{\mu(X_{t})}{\sigma(X_{t})}dt$ into the expanison (39) from Step 3, we have
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)$
$\qquad+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)du$
$\qquad+\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}dr\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)$
$\qquad+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)$
@@ Appendix C Proof of Step 4
$\qquad+\left((\tilde{\mu}^{\prime}(X_{s}))^{2}\sigma(X_{s})+\tilde{\mu}^{\prime\prime}(X_{s})\tilde{\mu}(X_{s})\sigma(X_{s})\right)$
$\qquad\qquad\times\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ\left(dB_{v}^{\mathbb{Q}}-\frac{\mu(X_{v})}{\sigma(X_{v})}dv\right)drdu$
$\qquad+\left(\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})+\tilde{\mu}^{\prime\prime}(X_{s})\tilde{\mu}(X_{s})\sigma(X_{s})\right)$
$\qquad\qquad\times\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}dv\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)du$
$\qquad+\left(\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})+\sigma^{\prime\prime}(X_{s})(\tilde{\mu}(X_{s}))^{2}\right)$
$\qquad\qquad\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}dvdr\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)$
$\qquad+\left(\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\sigma(X_{s})+\tilde{\mu}^{\prime\prime}(X_{s})(\sigma(X_{s}))^{2}\right)$
$\qquad\qquad\times\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ\left(dB_{v}^{\mathbb{Q}}-\frac{\mu(X_{v})}{\sigma(X_{v})}dv\right)\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)du$
$\qquad+\left(\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\sigma(X_{s})+\sigma^{\prime\prime}(X_{s})\sigma(X_{s})\tilde{\mu}(X_{s})\right)$
$\qquad\qquad\times\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ\left(dB_{v}^{\mathbb{Q}}-\frac{\mu(X_{v})}{\sigma(X_{v})}dv\right)dr\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)$
$\qquad+\left((\sigma^{\prime}(X_{s}))^{2}\tilde{\mu}(X_{s})+\sigma^{\prime\prime}(X_{s})\sigma(X_{s})\tilde{\mu}(X_{s})\right)$
$\qquad\qquad\times\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}dv\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)$
@@
$\qquad\times\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ\left(dB_{v}^{\mathbb{Q}}-\frac{\mu(X_{v})}{\sigma(X_{v})}dv\right)\circ\left(dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr\right)\circ\left(dB_{u}^{\mathbb{Q}}-\frac{\mu(X_{u})}{\sigma(X_{u})}du\right)$
@@
$X_{t}$ $=X_{s}-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{Q}}$
$\qquad+\left(\mu^{\prime}(X_{s})\sigma(X_{s})-\mu(X_{s})\sigma^{\prime}(X_{s})-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})^{2}-\frac{1}{2}\sigma(X_{s})^{2}\sigma^{\prime\prime}(X_{s})\right)$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\times\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{Q}}du$
$\qquad-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})^{2}\int_{0}^{t}\int_{0}^{s}du\circ dB_{s}^{\mathbb{Q}}+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}$
$+\left((\sigma^{\prime}(X_{s}))^{2}\sigma(X_{s})+\sigma^{\prime\prime}(X_{s})(\sigma(X_{s}))^{2}\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dB_{v}^{\mathbb{Q}}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}$
$\qquad+\varepsilon^{\tilde{\mu},\sigma}+\varepsilon^{\mu/\sigma}.$
@@
Appendix D Proof of Step 5
Comparing the coefficients of (20) from Step 2 and (23) from Step 4, for the expansion of $X_{t}$ at the initial point $X_{s}$ on interval $[s,t)$, we have the following equalities,
@@ Appendix D Proof of Step 5
Following the above observation, the parameter $\psi^{t_{i},t_{i+1}}_{21}$ can be found corresponding to the initial value $X_{t_{i}},\text{ for }i=0,1,2,\cdots,n$. With this, the equation (42) holds true for each $t_{i},i=0,1,2,\dots,n$. For example, (42) is equivalently written as
$\mu^{\prime}(X_{t_{i}})\sigma(X_{t_{i}})-\mu(X_{t_{i}})\sigma^{\prime}(X_{t_{i}})-\frac{1}{2}\sigma(X_{t_{i}})\sigma^{\prime}(X_{t_{i}})^{2}$- $\frac{1}{2}\sigma(X_{t_{i}})^{2}\sigma^{\prime\prime}(X_{t_{i}})=\psi_{21}^{t_{i},t_{i+1}},$ (43)
$\qquad\forall\;i=0,1,2,\dots,n.$